Find the particular integral of
(x-1)^2 y" - 2(x-1)y' + 2y = (x-1)^2 y(0) = 3, y'(0) = 6
1
Expert's answer
2020-07-20T17:15:14-0400

Given D.E is

"(x-1)^2 y'' - 2(x-1)y' + 2y = (x-1)^2 \:,y(0) = 3, y'(0) = 6"

Notice that solution of homogeneious equation "(x-1)^2 y'' - 2(x-1)y' + 2y =0" is

"y_1=(x-1)"

We find second independent solution fron Liouville formula is

"W_{y_1y_2}(x)=\begin{vmatrix}
 y_1&y_2\\y_1'&y_2'
\end{vmatrix}=C_1e^{-\int \frac{-2(x-1)}{(x-1)^2}dx}=C_1(x-1)^2"

Thus,

"\dfrac{y_1y_2'-y_1'y_2}{y_1^2}=C_1\frac{(x-1)^2}{(x-1)^2}=C_1\implies (\frac{y_2}{y_1})'=C_1\\
\implies y_2=C_1(x-1)^2+C_2(x-1)"

Hence, general solution of homogeneous equation is

"y_o(x)=C_1(x-1)^2+C_2(x-1)"

Now, using methood of variation of parameters and construct general solution of inhomogeneous equation ,

"y_1
(
x
)
 \:\text{and}\:
y_
2
(
x
)
:
y_
0
(
x
)
=
C_
1
y_
1
(
x
)
+
C_
2
y_
2
(
x
)
,
where \:
C_
1
,
C_
2
 \:\text{are arbitrary constants.}"

Thus,

"\text{The derivatives of the unknown functions }
C_
1
(
x
)
 \&
C_
2
(
x
)\:
\\ \text{can be determined from the system of equations}"


"C^
′_
1
(
x
)
y_
1
(
x
)
+
C^
′_
2
(
x
)
y_
2
(
x
)
=
0\\
C^
′_
1
(
x
)
y^
′_
1
(
x
)
+
C^
′_
2
(
x
)
y^
′_
2
(
x
)
=
f
(
x
)"

Hence,

"{{C’_1}\left( x \right) = – \frac{{{y_2}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}},\;\;}\kern-0.3pt
{{C’_2}\left( x \right) = \frac{{{y_1}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}.}"

Furthermore, knowing the derivatives of "{C’_1}\left( x \right)" and "{C’_2}\left( x \right)" ,one can find the "C_1(x)\& C_2(x)" Thus,

"{{{C_1}\left( x \right) }={ – \int {\frac{{{y_2}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} }+{ {A_1},\;\;}}\kern-0.3pt {{{C_2}\left( x \right) }={ \int {\frac{{{y_1}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} }+{ {A_2}}}"

where"A_1\&A_2" are constants of integration.Then the general solution of the original nonhomogeneous equation will be expressed by the formula


"{y\left( x \right) }={ {C_1}\left( x \right){y_1}\left( x \right) + {C_2}\left( x \right){y_2}\left( x \right) }
= {{\left[ { – \int {\frac{{{y_2}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} + {A_1}} \right] \cdot}\kern0pt{ {y_1}\left( x \right) }}
+\\ {{\left[ {\int {\frac{{{y_1}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} + {A_2}} \right] \cdot}\kern0pt{ {y_2}\left( x \right) }}
= {{{A_1}{y_1}\left( x \right) + {A_2}{y_2}\left( x \right) }+{ Y\left( x \right)}}"

Where,

"{Y\left( x \right) }
= {{y_2}\left( x \right)\int {\frac{{{y_1}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} }
– {{y_1}\left( x \right)\int {\frac{{{y_2}\left( x \right)f\left( x \right)}}{{{W_{{y_1},{y_2}}}\left( x \right)}}dx} }"

Thus, from above fact by putting "f(x)=(x-1)^2" and solving exactly as above we get the general solution is

"C_1(x-1)-(x+\ln(x-1))(x-1)+C_2x(x-1)+x\ln(x-1)(x-1)"

Thus particular solution is

"- (x-1)(10x+\ln(x-1)-x\ln(x - 1)+3-\pi i + \pi xi)"